# Crossover Designs in Software Engineering Experiments: Review of the State of Analysis

### v1

@@ Abstract.
Experimentation is an essential method for causal inference in any empirical discipline. Crossover-design experiments are common in Software Engineering (SE) research. In these, subjects apply more than one treatment in different orders. This design increases the amount of obtained data and deals with subject variability but introduces threats to internal validity like the learning and carryover effect. Vegas et al. reviewed the state of practice for crossover designs in SE research and provided guidelines on how to address its threats during data analysis while still harnessing its benefits. In this paper, we reflect on the impact of these guidelines and review the state of analysis of crossover design experiments in SE publications between 2015 and March 2024. To this end, by conducting a forward snowballing of the guidelines, we survey 136 publications reporting 67 crossover-design experiments and evaluate their data analysis against the provided guidelines. The results show that the validity of data analyses has improved compared to the original state of analysis. Still, despite the explicit guidelines, only 29.5% of all threats to validity were addressed properly. While the maturation and the optimal sequence threats are properly addressed in 35.8% and 38.8% of all studies in our sample respectively, the carryover threat is only modeled in about 3% of the observed cases. The lack of adherence to the analysis guidelines threatens the validity of the conclusions drawn from crossover design experiments.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: International Symposium on Empirical Software Engineering and Measurement; October 20–25, 2024; Barcelona, ESisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Empirical studiesccs: General and reference Experimentationccs: General and reference Design
@@ 1. Introduction
    We show gaps in SE literature on analyzing data from crossover design experiments.
@@ 2.1. Experimental Design
An important decision when designing an experiment is whether each subject is administered only one or multiple treatments. The former is referred to as an independent measure (or between-subject) design, where subjects are split into groups and each group applies only one treatment, the latter is referred to as a repeated measures (or within-subject) design. The benefit of a repeated measures design is two-fold. Firstly, the same amount of experimental subjects produce more data points than in an independent measure design. Secondly, comparing relative rather than absolute values of each response variable deals with subject variability. On the other hand, such a design introduces new threats to validity. Vegas et al. summarized four types of threats to validity in their guidelines (Vegas et al., 2015):
@@
    Subject variability: The tasks performed in SE experiments are often strongly influenced by differences between human subjects (Pickard et al., 1998) that are difficult to quantify.
  4. (4)
@@
The crossover design, a special form of the repeated-measures design where participants receive the treatments in different sequences (Vegas et al., 2015), counterbalances the threats by dispersing their impact on the response variable evenly. Still, the threats affect the observations of the response variable but are largely ignored in most SE papers (Vegas et al., 2015). In medical research, the carryover effect is often addressed via a washout period (Woods et al., 1989), i.e., a period between the experimental periods long enough for the medical compound to exit the subject’s system. In SE research, a washout period is often not feasible since it would require participants to unlearn techniques or tools they were subjected to (Kitchenham et al., 2003). Still, many SE papers even fail to acknowledge the carryover threat (Vegas et al., 2015).
Consequently, Vegas et al. recommended abandoning simple NHSTs that only test whether different treatments change the distribution of the response variable and rather adopting a linear mixed model (LMM) to analyze the data from a crossover-design experiment. An LMM can involve additional factors with an influence on the response variable, which (1) isolates the true effect of the main factor in question and (2) provides insight into whether the threats to validity apply in a particular instance of an experiment.
Figure 1 visualizes an LMM (second row) that determines the effect of treatment $t_{i}$ on a response variable $y_{i}$ while also addressing the above-mentioned threats (first row) by modeling factors that address those threats (third row). The bottom row visualizes which parts of the AB/BA crossover design experiment these terms affect. For example, the treatment $\beta_{t}t_{i}$ affects period 2 in sequence 1 and period 1 in sequence 2 (marked yellow in Figure 1). For simplicity, we constrain the visualization to a crossover design experiment with one main factor containing two levels (A and B), which requires two periods and two sequences (AB and BA) and is commonly referred to as AB/BA crossover design.
@@ 2.2. State of Practice
Since the introduction of evidence-based SE by Kitchenham et al. (Kitchenham et al., 2004), the field has been subject to reviews about the state of practice of various aspects of empirical research. For example, Kampenes et al. reviewed the state of practice of designing, conducting, and evaluating quasi-experiments (Kampenes et al., 2009). They conclude terminological ambiguity and a common lack of awareness of specific biases affect the validity of drawn conclusions. Menzies et al. reviewed data analysis practices in empirical SE research (Menzies and Shepperd, 2019) and consolidated 12 “bad smells” commonly committed in publications. Hannay et al. surveyed SE experiments regarding the degree of realism of the employed experimental material and task (Hannay and Jørgensen, 2008) and detected a lack of awareness of the implications of realism in SE publications.
The aforementioned work by Vegas et al. also contained a secondary study of crossover-design experiments (Vegas et al., 2015), which motivated the guidelines described in Section 2.1. Their review identified similar problems in the state of practice. Several publications misuse terminology, remain unclear in their design decision and apply an incorrect analysis to the data from crossover-design experiments. These results are supported by a similar study by Kitchenham et al. which focused on families of experiments (Kitchenham et al., 2019).
@@ 3.1. Study Inclusion
The first and second authors of this study conducted the inclusion phase. The 136 articles were distributed among the two authors based on their availability (90 for the first author, 46 for the second). For each paper, the assigned author read the abstract, introduction, and method section to decide the inclusion and exclusion criteria. Only criterion Ex3 (i.e., filtering out duplicates or extensions) was performed centrally by the first author by clustering the sample by author names and investigating similar candidates. During the inclusion phase, unclear decisions were flagged and later reviewed by the third author of this study, a senior researcher and expert in the field of experimentation methodology. The third author acted as an arbiter and decided on the three unclear cases. The inclusion phase identified 48 primary studies ($48/136=35.3\%$) as eligible for the subsequent data extraction phase.
@@ 3.2. Data Extraction
Location URL of the material if available Text
@@
We summarized the extraction guidelines containing a definition, concise extraction rules, as well as examples in a shared document. To assess the reliability of these guidelines, the first and second authors performed an overlap of the extraction task prior to the main extraction phase. During this overlap, the two authors applied the extraction guidelines to the seven primary studies from the 14 that were already involved in the inclusion overlap and which were included according to our criteria. We assessed the agreement of all attributes from the subjects, analysis, and material group except the location attribute as it bears no empirical value. We calculated Pearson’s correlation coefficient (PCC) for the numerical attribute subject number and Bennett’s S-score (Bennett et al., 1954) between the two ratings of categorical attributes. None of the seven primary studies in our subset applied a washout period, which is why we did not calculate the inter-rater agreement for this variable. We selected Bennett’s S-score over the more common Cohen’s Kappa as the latter is known to be unreliable for uneven marginal distributions (Feng, 2015). The two raters were in perfect agreement about the subject number attribute ($PCC=1.0$) and achieved an average S-score of 85.8% over all categorical columns. After confirming the reliability of the extraction guidelines, the two first authors proceeded with applying them to the 48 primary studies originally assigned to them during the inclusion phase. The third author acted as an arbiter and clarified seven unclear instances in the data extraction.
@@ 3.3. Data Analysis
Upon completion of the data extraction phase, we generated descriptive statistics from the distribution of attribute values. We visualized categorical data using bar charts and numerical data using box plots. The main attribute of interest, the threat addressal, was represented as a heatmap where one axis listed the four threats to validity and the other axis the types of threat addressal.
@@ 4. Results
The 48 primary studies describe 67 experiments. Main factors of interest include, among others, testing paradigms, like test-driven development (Baldassarre et al., 2022; Esposito et al., 2023; Fucci et al., 2016), the effect of API design rules (Bogner et al., 2023), and noise during software development (Romano et al., 2020). The following subsections report the obtained results grouped by the attribute groups in Table 2.
@@ 4.3. Material
Figures 7 and 8 visualize the availability of data sets and analysis script. [^2] Note that these are distributions among the 48 publications, not among the 67 experiments, because, in our sample, we observed that any additional material was always associated with a publication, not an individual experiment. The majority of material (20 data sets and 30 analysis scripts are unavailable) were never available and a portion (4 data sets and 4 analysis scripts are broken) has become unavailable since their original publication. Among the remaining material, several data sets ($12/48=25\%$) and scripts ($12/48=22.9\%$) have been properly archived and, therefore, preserved for future use like reproduction.
@@ 6.1. Threats to Validity
Our study suffers from the following threats to validity (Wohlin et al., 2012). Most prominently, our study is subject to a threat to external validity. While we aim to draw general conclusions about experimentation in SE literature, our sample is limited to SE literature that cites the investigated guidelines (Vegas et al., 2015). However, we argue that this sample is adequate for this paper for the following reasons. The guidelines (Vegas et al., 2015) are—to the best of our knowledge—the only SE-specific guidelines for analyzing crossover-design experiments in SE. Hence, authors of our selected subset had access to the only guidelines explicitly advising them on how to properly analyze their experimental data. Therefore, we are confident that our sample represents an upper bound to extrapolate to the SE experimentation literature.
@@ 6.2. Future Work
In the scope of this study, we only surveyed how authors addressed the four threats to validity presented in Section 2.1 as per the original guidelines (Vegas et al., 2015). We did not consider additional threats to validity which crossover-design experiments are subject to. An important, additional threat to consider in future work is the material threat, i.e., the influence of the used experimental material (e.g., code snippet). In a $2\times 2$-crossover design experiment, the material threat is inherently confounded either with the period (if both sequence groups use the same material in each period) or the treatment (if both sequence groups use the same material for each level of the main factor). Blocking the threat exhibited by experimental material requires additional periods in the crossover design, which some papers in our sample did (Albuquerque et al., 2023; Kreber and Diehl, 2023; Sandobalin et al., 2020). Other additional threats to validity to consider in future evaluations of the state of practice include interaction effects between subjects and treatments (e.g., personal preference for specific treatments).
In future work, we plan to reproduce the analysis of the surveyed experiments as far as possible. While we cannot reproduce the analysis of experiments where the raw data is not available, we plan to (1) reproduce the original analyses as described in the publication, and (2) reanalyze the data according to the data analysis guidelines (Vegas et al., 2015). Contrasting these analyses will produce more examples for an application of the guidelines, but also reveal studies that drew incorrect conclusions due to incorrect data analyses. This quantifies the effect of the guidelines in terms of the risk of drawing inappropriate conclusions and will serve as further motivation for adherence. Additionally, we plan to analyze and reproduce crossover-design experiments that do not cite the data analysis guidelines to compare them with our current sample and, therefore, further study the guidelines’ impact.
@@ 7. Conclusion
In this reflection, we investigate the state of practice of analyzing crossover design experiments in SE. A sample of publications citing explicit analysis guidelines (Vegas et al., 2015) shows that the state of practice still contains several, significant gaps threatening the validity of the drawn conclusions. While the guidelines by Vegas et al. (Vegas et al., 2015) supported a significant portion of authors to analyze their obtained data at least in parts, the general sample of papers shows room for improvement.
We hope that the overview of the state of practice of analyzing crossover design experiments in SE encourages authors to investigate this topic more thoroughly and that both our visualizations as well as the recovery of the data analysis script from the original guidelines (Vegas et al., 2015) will support the latter in guiding authors towards a correct analysis. We encourage authors to abandon the overly simple NHST analysis (Frattini et al., 2024b) for more complex (G)LMMs, which enable them to adequately address threats to the validity of crossover-design experiments during analysis.

### v2

@@ Abstract.
Experimentation is an essential method for causal inference in any empirical discipline. Crossover-design experiments are common in Software Engineering (SE) research. In these, subjects apply more than one treatment in different orders. This design increases the amount of obtained data and deals with subject variability but introduces threats to internal validity like the learning and carryover effect. Vegas et al. reviewed the state of practice for crossover designs in SE research and provided guidelines on how to address its threats during data analysis while still harnessing its benefits. In this paper, we reflect on the impact of these guidelines and review the state of analysis of crossover-design experiments in SE publications between 2015 and March 2024. To this end, by conducting a forward snowballing of the guidelines, we survey 136 publications reporting 67 crossover-design experiments and evaluate their data analysis against the provided guidelines. The results show that the validity of data analyses has improved compared to the original state of analysis. Still, despite the explicit guidelines, only 29.5% of all threats to validity were addressed properly. While the maturation and the optimal sequence threats are properly addressed in 35.8% and 38.8% of all studies in our sample respectively, the carryover threat is only modeled in about 3% of the observed cases. The lack of adherence to the analysis guidelines threatens the validity of the conclusions drawn from crossover-design experiments.
@@
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3674805.3690754conference: International Symposium on Empirical Software Engineering and Measurement; October 20–25, 2024; Barcelona, ESisbn: 979-8-4007-1047-6/24/10ccs: General and reference Empirical studiesccs: General and reference Experimentationccs: General and reference Design
@@ 1. Introduction
    We show gaps in SE literature on analyzing data from crossover-design experiments.
@@ 2.1. Experimental Design
An important decision when designing an experiment is whether each subject is administered only one or multiple levels of the main factor. The former is referred to as an independent measures (or between-subjects) design, where subjects are split into groups and each group applies only one treatment, the latter is referred to as a repeated measures (or within-subjects) design. The benefit of a repeated measures design is two-fold. Firstly, the same amount of experimental subjects produce more data points than in an independent measure design. Secondly, comparing relative rather than absolute values of each response variable deals with the threat of subject variability, i.e., that tasks performed in SE experiments are often influenced by differences between human subjects (Pickard et al., 1998) which are difficult to quantify. On the other hand, such a design introduces new threats to validity. Vegas et al. identified three additional types of threats in their guidelines (Vegas et al., 2015):
@@
The crossover design, a special form of the repeated-measures design where participants receive the treatments in different sequences (Vegas et al., 2015), counterbalances the threats by dispersing their impact on the response variable evenly. Still, the threats (including the threat of subject variability) affect the observations of the response variable but are largely ignored in most SE papers (Vegas et al., 2015). In medical research, the carryover effect is often addressed via a washout period (Woods et al., 1989), i.e., a period between the experimental periods long enough for the medical compound to exit the subject’s system. In SE research, a washout period is often not feasible since it would require participants to unlearn techniques or tools they were subjected to (Kitchenham et al., 2003). Still, many SE papers even fail to acknowledge the carryover threat (Vegas et al., 2015).
Consequently, Vegas et al. recommended abandoning simple NHSTs that only test whether different treatments change the distribution of the response variable and rather adopting linear mixed models (LMMs) to analyze the data from crossover-design experiments. An LMM can involve additional factors with an influence on the response variable, which (1) isolates the true effect of the main factor in question and (2) provides insight into whether the threats to validity apply in a particular instance of an experiment.
Figure 1 visualizes an LMM (second row) that determines the effect of treatment $t_{i}$ on a response variable $y_{i}$ while also addressing the above-mentioned threats (first row) by modeling factors that address those threats (third row). The bottom row visualizes which parts of the AB/BA crossover-design experiment these terms affect. For example, the treatment $t_{i}$ affects period 2 in sequence 1 and period 1 in sequence 2 (marked yellow in Figure 1) with strength $\beta_{t}$. For simplicity, we constrain the visualization to a crossover-design experiment with one main factor containing two levels (A and B), which requires two periods and two sequences (AB and BA) and is commonly referred to as AB/BA crossover design.
@@ 2.2. State of Practice
Since the introduction of evidence-based SE by Kitchenham et al. (Kitchenham et al., 2004), the field has been subject to reviews about the state of practice of various aspects of empirical research. For example, Kampenes et al. reviewed the state of practice of designing, conducting, and evaluating quasi-experiments (Kampenes et al., 2009). They conclude that terminological ambiguity and a lack of awareness of specific biases affect the validity of drawn conclusions. Menzies et al. reviewed data analysis practices in empirical SE research (Menzies and Shepperd, 2019) and consolidated 12 “bad smells” commonly committed in publications. Hannay et al. surveyed SE experiments regarding the degree of realism of the employed experimental material and task (Hannay and Jørgensen, 2008) and detected a lack of awareness of the implications of realism in SE publications.
The aforementioned work by Vegas et al. also contained a secondary study of crossover-design experiments (Vegas et al., 2015), which motivated their guidelines described in Section 2.1. The review identified similar problems in the state of practice. Several publications misuse terminology, remain unclear in their design decision, and apply an incorrect analysis to the data from crossover-design experiments. These results are supported by a similar study by Kitchenham et al. which focused on families of experiments (Kitchenham et al., 2019).
@@ 3.1. Study Inclusion
The first and second authors of this study conducted the inclusion phase. The 136 articles were distributed among the two authors based on their availability (90 for the first author, 46 for the second). For each paper, the assigned author read the abstract, introduction, and method section to decide the inclusion and exclusion criteria. Only criterion Ex3 (i.e., filtering out duplicates or extensions) was performed centrally by the first author by clustering the sample by author names and investigating similar candidates. During the inclusion phase, unclear decisions were flagged and later reviewed by the third author of this study. The third author acted as an arbiter and decided on the three unclear cases. The inclusion phase identified 48 primary studies ($48/136=35.3\%$) as eligible for the subsequent data extraction phase.
@@ 3.2. Data Extraction
Location URL of the material if available URL
@@
We summarized the extraction guidelines containing a definition, concise extraction rules, as well as examples for each attribute in a shared document. To assess the reliability of these guidelines, the first and second authors performed an overlap of the extraction task prior to the main extraction phase. During this overlap, the two authors applied the extraction guidelines to the seven primary studies from the 14 that were already involved in the inclusion overlap and which were included according to our criteria. We assessed the agreement of all attributes from the subjects, analysis, and material group except the location attribute as it bears no empirical value. We calculated Pearson’s correlation coefficient (PCC) for the numerical attribute subject number and Bennett’s S-score (Bennett et al., 1954) between the two ratings of categorical attributes. None of the seven primary studies in our subset applied a washout period, which is why we did not calculate the inter-rater agreement for this variable. We selected Bennett’s S-score over the more common Cohen’s Kappa as the latter is known to be unreliable for uneven marginal distributions (Feng, 2015). The two raters were in perfect agreement about the subject number attribute ($PCC=1.0$) and achieved an average S-score of 85.8% over all categorical columns. After confirming the reliability of the extraction guidelines, the two first authors proceeded with applying them to the 48 primary studies originally assigned to them during the inclusion phase. The third author acted as an arbiter and clarified seven unclear instances in the data extraction.
@@ 3.3. Data Analysis
Upon completion of the data extraction phase, we generated descriptive statistics from the distribution of attribute values. We visualized categorical data using bar charts and numerical data using box plots. We represented the main attribute of interest, the threat addressal, as a heatmap where one axis listed the four threats to validity and the other axis the types of threat addressal.
@@ 4. Results
The 48 primary studies describe 67 experiments. Main factors of interest include, among others, testing paradigms like test-driven development (Baldassarre et al., 2022; Esposito et al., 2023; Fucci et al., 2016), the effect of API design rules (Bogner et al., 2023), and noise during software development (Romano et al., 2020). The following subsections report the obtained results grouped by the attribute groups in Table 2.
@@ 4.3. Material
Figures 7 and 8 visualize the availability of data sets and analysis script. [^2] Note that these are distributions among the 48 publications, not among the 67 experiments, because, in our sample, we observed that any additional material was always associated with a publication, not an individual experiment. Most of the material was never available (20 data sets and 30 scripts are unavailable) and a portion has become unavailable since their original publication (4 data sets and 4 scripts are broken). Among the remaining material, several data sets ($12/48=25\%$) and scripts ($12/48=22.9\%$) have been properly archived and, therefore, preserved for future use like reproduction (Frattini et al., 2024c).
@@ 5. Discussion
To improve future analyses of crossover-design experiments in SE, we recommend increasing the awareness of established guidelines (Vegas et al., 2015). Including these guidelines in textbooks (Wohlin et al., 2012) and standards (Ralph et al., 2020) will aid authors in more rigorous analyses. Furthermore, we urge reviewers to put more emphasis on guideline adherence. This not only requires awareness of the existence of guidelines (i.e., checking that appropriate guidelines were cited) but also of their content (i.e., checking that the guidelines were properly followed).
@@ 6.1. Threats to Validity
Our study suffers from the following threats to validity (Wohlin et al., 2012). Most prominently, our study is subject to a threat to external validity. While we aim to draw general conclusions about experimentation in SE literature, our sample is limited to SE literature that cites the investigated guidelines (Vegas et al., 2015). However, we argue that our sample is adequate for this paper for the following reasons. The guidelines (Vegas et al., 2015) are—to the best of our knowledge—the only SE-specific guidelines for analyzing crossover-design experiments in SE. Hence, authors of our selected subset had access to the only guidelines explicitly advising them on how to properly analyze their experimental data. Therefore, we are confident that our sample represents an upper bound to extrapolate to the SE experimentation literature.
@@ 6.2. Future Work
In the scope of this study, we only surveyed how authors addressed the four threats to validity presented in Section 2.1 as per the original guidelines (Vegas et al., 2015). We did not consider additional threats to validity which crossover-design experiments are subject to. Additional threats we plan to investigate is the material threat, i.e., the influence of the used experimental material (e.g., code snippet), and the interaction effects between subjects and treatments, e.g.,, personal preference for specific treatments.
Furthermore, we plan to extend our sample by including SE studies presenting experiments with a crossover design that do not cite the guidelines by Vegas et al. (Vegas et al., 2015). By comparing this sample with our current one we aim to further characterize the impact of these guidelines.
Finally, we plan to reproduce the analysis of the surveyed experiments as far as possible. While we cannot reproduce the analysis of experiments where the raw data is not available, we plan to (1) reproduce the original analyses as described in the publication, and (2) reanalyze the data according to the data analysis guidelines (Vegas et al., 2015). Contrasting these analyses will produce more examples for an application of the guidelines, but also reveal studies that drew incorrect conclusions due to incorrect data analyses. This quantifies the effect of the guidelines in terms of the risk of drawing inappropriate conclusions and will serve as further motivation for adherence. Additionally, we plan to analyze and reproduce crossover-design experiments that do not cite the data analysis guidelines to compare them with our current sample and, therefore, further study the guidelines’ impact.
@@ 7. Conclusion
In this reflection, we investigate the state of practice of analyzing crossover-design experiments in SE. A sample of publications citing explicit guidelines (Vegas et al., 2015) shows that the state of practice still contains several, significant gaps threatening the validity of the drawn conclusions. While the guidelines by Vegas et al. (Vegas et al., 2015) supported a significant portion of authors to analyze their obtained data at least in parts, the general sample of papers shows room for improvement.
We hope that the overview of the state of practice of analyzing crossover-design experiments in SE encourages authors to investigate this topic more thoroughly and that both our visualizations as well as the recovery of the data analysis script from the original guidelines (Vegas et al., 2015) will support the latter in guiding authors towards a correct analysis. We encourage authors to abandon the overly simple NHST analysis (Frattini et al., 2024b) for more complex (G)LMMs, which enable them to adequately address threats to the validity of crossover-design experiments during analysis.